OpenStudy (anonymous):

Please help. What does it mean? Write in standard form. 10 to the 3 power.

OpenStudy (anonymous):

i guess it means \(10^3=10\times 10\times 10=1000\)

OpenStudy (anonymous):

oh ok thanks